# Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy Lensing

### v1

@@ Abstract
[…]
@@
pubyear: 2015pagerange: Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy LensingB
@@ 1 Introduction
With large imaging surveys like the Dark Energy Survey (DES) (Abbott et al., 2022), the Kilo-Degree Survey (KiDS) (Heymans et al., 2021), and the Hyper Suprime-Cam Survey (HSC) (Sugiyama et al., 2023), we have seen a rapid change in our ability to probe the large scale structure (LSS) and matter distribution in the non-linear regime using weak lensing (WL) and clustering. One of the biggest challenges standing in the way of using such measurements to test the dark components of the standard model is the modelling of baryons. In particular, baryonic feedback suppresses the matter power spectrum on non-linear scales, and therefore the weak lensing amplitude, and must be modelled accurately to avoid biasing cosmological inference (e.g., van Daalen et al., 2011a; Chisari et al., 2019a; Bigwood et al., 2024).
However, the complexity of baryonic feedback requires a large dynamic range in resolution and time that is difficult to simulate (e.g. Crain & van de Voort, 2023). This has led to the development of simplified ‘subgrid’ prescription models in hydrodynamical cosmological simulations which are used to mimic the effects of baryonic feedback on LSS (Steinborn et al., 2015; Chisari et al., 2019a; Schaye et al., 2023; Pakmor et al., 2023). Although the simulations are able to reproduce a range of observations, they give a wide spread of predictions for the amplitude and scale extent of the suppression of the matter power spectrum due to baryonic feedback (e.g., Chisari et al., 2019a; Villaescusa-Navarro et al., 2021; Sunseri et al., 2023). Understanding and reducing this model uncertainty is the challenge. More recently, it has been possible to constrain the matter power suppression from a number of observables (e.g. Debackere et al., 2020; Grandis et al., 2024; Bigwood et al., 2024; La Posta et al., 2024; Medlock et al., 2025), although these constraints are still not very precise.
@@
The kinetic Sunyaev-Zeldovich (kSZ) effect is a secondary anisotropy in the CMB that comes from the Doppler boost of CMB photons by inverse Compton scattering of free electrons in the Intracluster Medium (ICM) and Circumgalactic Medium (CGM) that have bulk motion along the line of sight relative to the CMB rest frame (Zeldovich & Sunyaev, 1969; Mroczkowski et al., 2019). The kSZ effect is small relative to other CMB secondaries (e.g. tSZ), follows a blackbody spectrum, and is directly proportional to the product of the density of electrons surrounding galaxies and their bulk velocity along the line of sight. The dependency on the bulk line of sight velocity and the small amplitude relative to other secondaries requires methods such as a velocity weighted stack using external velocity measurements from galaxy surveys to measure the signal. Stacked measurements with a velocity reconstruction weighting have revealed that the electrons around galaxies in the Baryon Oscillation Spectroscopic Survey (BOSS) are significantly more extended than the dark matter distribution (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a). The kSZ effect is particularly useful for WL as it probes the density of electrons (and therefore baryons) in the outskirts of the halo ($\sim$1-4 virial radii) which lies in the range of scales that baryons need to be better understood to improve WL cosmological constraints. Joint analyses of the kSZ effect with WL measurements are a promising way to constrain baryonic feedback and simultaneously improve cosmological constraints (e.g., Bigwood et al., 2024).
@@
In order to accurately interpret the kSZ effect, one must know the halo mass of the sample (Moser et al., 2021) and have an appropriate forward model as was done in Amodeo et al. (2021). By jointly analysing the kSZ effect with GGL measurements we can build a self-consistent framework for estimating the halo mass. Without GGL, analyses of the kSZ effect either rely on estimates of stellar mass (Maraston & Strömbäck, 2011; Maraston et al., 2013) which are then used to estimate the halo mass through stellar-to-halo mass relations (Miyatake et al., 2015; Kravtsov et al., 2018), or fits to galaxy clustering (Miyatake et al., 2015; Rodríguez-Torres et al., 2016), both of which have greater uncertainties than galaxy-galaxy lensing estimates. Bigwood et al. (2024) estimated that for the BOSS CMASS galaxy sample, these estimates span almost an order of magnitude. In this work we present the first joint fits of kSZ measurements and GGL measurements to disentangle baryons from dark matter in the outskirts of galactic halos, and more accurately constrain the mass of the sample. Accurate mass estimates where the baryon content is considered is also important for cluster cosmology (e.g. Cromer et al., 2022).
@@
Measurements of the kSZ effect can be used to test baryon feedback models and hydrodynamical simulations. In this case, having an accurate halo mass of the sample is as important, in order to interpret the appropriate strength of feedback (McCarthy et al., 2024). Previous analyses of the BOSS kSZ measurements of Schaan et al. (2021) are consistent with feedback that is stronger than most hydrodynamical simulations predict (Bigwood et al., 2024; McCarthy et al., 2024). This underestimation may be severe enough that it may even bias WL cosmological constraints (Amon & Efstathiou, 2022; Preston et al., 2023). In this work we build a fast analytic modelling framework based on the halo model, complementary to previous approaches that rely on the baryonification method, or expensive hydrodynamical simulations. This framework can be used for cosmological analyses of GGL and clustering in the future.
@@
In this work, we build a model to consistently describe the excess surface mass density of halos, as measured by GGL (Amon et al., 2023), with the gas distribution, probed by measurements of the kSZ effect (Schaan et al., 2021). We describe these measurements in more detail below.
@@ 2.1 Galaxy-Galaxy Lensing
The KiDS-1000 weak lensing data were taken with the 2.6 m VLT Survey Telescope (VST) (Capaccioli & Schipani, 2011; Grado et al., 2012) and span 777 $\rm deg^{2}$ of the sky (Giblin et al., 2021) with a number density of 6.22 $\rm arcmin^{-2}$.
@@
The kSZ effect measurements used in this work were obtained by Schaan et al. (2021) by stacking the CMB temperature fluctuations mapped by ACT (in 98 GHz and 150 GHz frequencies) around BOSS CMASS galaxies (DR10), weighted by their peculiar velocities. The covariance matrix is constructed using the bootstrap method. [^2] Higher signal-to-noise kSZ measurements have recently been reported in (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a) using ACT DR6 data; our method could be extended in future to incorporate these data.
@@
The measurements span a range of aperture radii $\theta_{\rm d}$ from 1 to 6 arcminutes which probes the kSZ profile in the range of 0.5 to 4 virial radii in the CMASS galaxy sample. These scales were chosen to probe baryonic feedback while minimizing covariance in the measurements. Schaan et al. (2021) stacked the measured temperatures $\mathcal{T}_{i}(\theta_{\rm d})$ around each galaxy in the sample. The stacking is done using a velocity weighted, inverse-variance weighted mean estimator given by
@@
where $\mathbf{v}$ is the velocity field, $f=d\ln\delta/d\ln a$ is the logarithmic linear growth rate, $\mathbf{\hat{n}}$ is the line-of-sight, $a$ is the scale factor, $H$ is the Hubble parameter, $b$ is the linear bias, and $\delta_{\rm g}$ is the galaxy overdensity. Solving the linearized continuity equation allows for the reconstruction of the radial velocity field which can be assigned to each galaxy in the sample. The estimator is carefully chosen so that the radial direction of the velocity field does not cancel out thus leaving the kSZ signal intact. Lastly, the measurements of Schaan et al. (2021) perform a bootstrap resampling of the individual galaxies to estimate the covariance of the stacked kSZ profiles. The measurements at smaller aperture sizes have a dominant source of noise coming from the detector and foregrounds while the measurements at larger aperture sizes are dominated by large-scale CMB fluctuations.
@@ 3 Halo model
The first term refers to the 1-halo density profile which corresponds to density from the effective central halo, and the second term refers to the 2-halo density profile which corresponds to contributions in the density coming from the surrounding neighbouring halos (Asgari et al., 2023). We model the transition between the 1-halo and 2-halo terms as a sum. This simplistic treatment of the transition region should suffice for the current level of precision in our analysis. In the subsequent subsections below we will explain these 1-halo and 2-halo terms in more detail.
We use a fixed Planck flat $\Lambda$CDM cosmology with the following values for parameters: $\Omega_{\rm m}=0.3111$, $\Omega_{\rm b}=0.0490$, $h=0.6766$, $n_{s}=0.9665$, $\sigma_{8}=0.8102$ (Planck Collaboration et al., 2020b). We fix the redshift of our model to be $\langle z\rangle=0.55$, the mean redshift of our lens sample (BOSS CMASS). This choice does not impact the results of our analysis when compared to a marginalization over the redshift distribution of the lens sample as shown in Amodeo et al. (2021). We further verified this to be true while including GGL.
@@ 3.1 The One Halo Term
The 1-halo term of the density profile is the contribution to the density from the central halo itself. For this work we assume the host galaxy lies at the centre of a dark matter halo and that the gas is bound to the dark matter halo (i.e. no miscentering). In other words, we do not model the gas and dark matter around satellites and centrals with a Halo-Occupation Distribution (HOD) framework. This model ignores the potential systematic effect of satellites or ‘miscentering’, where satellite galaxies offset the stacking procedures used on the central galaxies in both the kSZ effect measurements and the GGL measurements.
We also ignore an additional uncertainty on the reconstruction bias factor $r_{v}$ which is about $\sim 10\%$ for the reconstruction used in this analysis (Schaan et al., 2021). We note that modelling of more precise measurements should include this effect (Ried Guachalla et al., 2024; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025a).
@@ 3.1.1 Dark Matter
to determine the concentration given a halo mass $c(M_{\rm halo})$ at $z=0.55$. By using $\Delta=200$, our definition of halo mass $M_{\rm halo}$ is the same as $M_{200}$. This concentration relation was determined from N-body simulations for halo masses $10^{11}-10^{15}\;\rm M_{\odot}$ and redshifts $20<z<$ 2 which matches our selected sample. We use a comoving $\bar{\rho}_{\rm m}$ because the radius $r$ is in comoving units. With this
@@
In this work, we don’t model the potential effects of assembly bias from the surrounding environment in our CMASS sample. This effect could bias our results at the few percent level (Yuan et al., 2021; Amon et al., 2023) and is a necessary improvement for future work. We also choose to ignore the back-reaction of baryonic feedback on the dark matter distribution: a second-order effect which is not necessary to model for our current level of precision (van Daalen et al., 2011b; Springel et al., 2018). Our model of the 1-halo term closely follows an approach from Cromer et al. (2022) which modelled baryons in WL of clusters. This approach differs from HOD based halo models commonly used to model GGL (Nishimichi et al., 2019; Miyatake et al., 2022; Zacharegkas et al., 2022; Dvornik et al., 2023), but it does allow for an explicit modelling of the baryon density profile which can be used to jointly analyse the kSZ measurements enabling constraints on baryon density parameters and dark matter parameters simultaneously.
@@
$x_{c}$ Core scale fraction $\mathcal{U}(0.0,1.0)$ $>0.48$ $>0.46$ $>0.49$
@@
$\beta$ Outer slope $\mathcal{U}(1,5)$ $2.47^{+0.34}_{-0.46}$ $<3.51$ $2.45^{+0.30}_{-0.23}$
@@
$A_{\rm 2h}$ Two halo term amplitude $\mathcal{U}(0.0,1.5)$ $>0.631$ $0.66\pm 0.16$ $0.64\pm 0.16$

### v2

@@ Abstract
We present the first joint analysis of the kinetic Sunyaev-Zeldovich (kSZ) effect with galaxy-galaxy lensing (GGL) for CMASS galaxies in the Baryon Oscillation Spectroscopic Survey (BOSS). We show these complementary probes can disentangle baryons from dark matter in the outskirts of galactic halos by alleviating model degeneracies that are present when fitting to kSZ or GGL measurements alone. In our joint kSZ+GGL analysis we show that the baryon density profile is well constrained on scales from 0.3 to 50 Mpc/$h$. With our well constrained profile of the baryon density, we provide direct comparisons to simulations. For our model we find an outer slope of the baryon distribution that is shallower than predicted by some hydrodynamical simulations, consistent with enhanced baryonic feedback. We also show that not including baryons in a model for GGL can shift halo mass estimates by $\sim 20\%$ compared to a model that includes baryons and is jointly fit to kSZ+GGL measurements. Our modelling code galaxy-galaxy lensing and kSZ (glasz) is publicly available at https://github.com/James11222/glasz.
@@
pubyear: 2026pagerange: Disentangling the Halo: Joint Model for Measurements of the Kinetic Sunyaev-Zeldovich Effect and Galaxy-Galaxy LensingC
@@ 1 Introduction
With large imaging surveys like the Dark Energy Survey (DES) (Abbott et al., 2022), the Kilo-Degree Survey (KiDS) (Heymans et al., 2021), and the Hyper Suprime-Cam Survey (HSC) (Sugiyama et al., 2023), we have seen a rapid change in our ability to probe the large scale structure (LSS) and matter distribution in the non-linear regime using weak lensing (WL) and clustering. One of the biggest challenges standing in the way of using such measurements to test the dark components of the standard model is the modelling of baryons. In particular, baryonic feedback suppresses the matter power spectrum on non-linear scales, and therefore the weak lensing amplitude, and must be modelled accurately to avoid biasing cosmological inference (e.g., van Daalen et al., 2011b; Chisari et al., 2019b; Bigwood et al., 2024).
However, the complexity of baryonic feedback requires a large dynamic range in resolution and time that is difficult to simulate (e.g. Crain and van de Voort, 2023). This has led to the development of simplified ‘subgrid’ prescription models in hydrodynamical cosmological simulations which are used to mimic the effects of baryonic feedback on LSS (Steinborn et al., 2015; Chisari et al., 2019b; Schaye et al., 2023; Pakmor et al., 2023). Although the simulations are able to reproduce a range of observations, they give a wide spread of predictions for the amplitude and scale extent of the suppression of the matter power spectrum due to baryonic feedback (e.g., Chisari et al., 2019b; Villaescusa-Navarro et al., 2021; Sunseri et al., 2023). Understanding and reducing this model uncertainty is the challenge. More recently, it has been possible to constrain the matter power suppression from a number of observables (e.g. Debackere et al., 2020; Grandis et al., 2024; Bigwood et al., 2024; La Posta et al., 2024; Medlock et al., 2025), although these constraints are still not very precise.
@@
The kinetic Sunyaev-Zeldovich (kSZ) effect is a secondary anisotropy in the CMB that comes from the Doppler boost of CMB photons by inverse Compton scattering of free electrons in the Intracluster Medium (ICM) and Circumgalactic Medium (CGM) that have bulk motion along the line of sight relative to the CMB rest frame (Zeldovich and Sunyaev, 1969; Mroczkowski et al., 2019). The kSZ effect is small relative to other CMB secondaries (e.g. tSZ), follows a blackbody spectrum, and is directly proportional to the product of the density of electrons surrounding galaxies and their bulk velocity along the line of sight. The dependency on the bulk line of sight velocity and the small amplitude relative to other secondaries requires methods such as a velocity weighted stack using external velocity measurements from galaxy surveys to measure the signal. Stacked CMB measurements from the Atacama Cosmology Telescope (ACT) survey with a velocity reconstruction weighting have revealed that the electrons around galaxies in the Baryon Oscillation Spectroscopic Survey (BOSS) are significantly more extended than the dark matter distribution (Schaan et al., 2016, 2021; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b). The kSZ effect is particularly useful for WL as it probes the density of electrons (and therefore baryons) in the outskirts of the halo ($\sim$1-4 virial radii) which lies in the range of scales that baryons need to be better understood to improve WL cosmological constraints. Joint analyses of the kSZ effect with WL measurements are a promising way to constrain baryonic feedback and simultaneously improve cosmological constraints (e.g., Bigwood et al., 2024).
@@
In order to accurately interpret the kSZ effect, one must know the halo mass of the sample (Moser et al., 2021) and have an appropriate forward model as was done in Amodeo et al. (2021). By jointly analysing the kSZ effect with GGL measurements we can build a self-consistent framework for estimating the halo mass. Without GGL, analyses of the kSZ effect either rely on estimates of stellar mass (Maraston and Strömbäck, 2011; Maraston et al., 2013) which are then used to estimate the halo mass through stellar-to-halo mass relations (Miyatake et al., 2015; Kravtsov et al., 2018), or fits to galaxy clustering (Miyatake et al., 2015; Rodríguez-Torres et al., 2016), both of which have greater uncertainties than galaxy-galaxy lensing estimates. Bigwood et al. (2024) estimated that for the BOSS CMASS galaxy sample, these estimates span almost an order of magnitude. In this work we present the first joint fits of kSZ measurements and GGL measurements to disentangle baryons from dark matter in the outskirts of galactic halos, and more accurately constrain the mass of the sample. Accurate mass estimates where the baryon content is considered is also important for cluster cosmology (e.g. Cromer et al., 2022).
@@
Measurements of the kSZ effect can be used to test baryon feedback models and hydrodynamical simulations. In this case, having an accurate halo mass of the sample is as important, in order to interpret the appropriate strength of feedback (McCarthy et al., 2024). Previous analyses of the BOSS kSZ measurements of Schaan et al. (2021) are consistent with feedback that is stronger than most hydrodynamical simulations predict (Bigwood et al., 2024; McCarthy et al., 2024). This underestimation may be severe enough that it may even bias WL cosmological constraints (Amon and Efstathiou, 2022; Preston et al., 2023). In this work we build a fast analytic modelling framework based on the halo model, complementary to previous approaches that rely on the baryonification method, or expensive hydrodynamical simulations. While out of the scope of this work, validation of our model with simulations will be necessary to use this model for cosmological analyses of GGL and clustering in the future.
@@
In this work, we build a model to consistently describe the excess surface mass density of halos, as measured by GGL (Amon et al., 2023), with the gas distribution, probed by measurements of the kSZ effect (Schaan et al., 2021). These measurements require data from a combination of several surveys (BOSS, DES Y3, KiDS-1000, ACT DR5), for clarity we summarize the sky footprints and redshift distributions of these surveys in Figure 1. We describe these measurements in more detail below.
Figure 1: Left: Sky footprints in equatorial coordinates of all surveys used in this work. Spectroscopic lenses are from the BOSS CMASS (purple) survey, photometric sources used for weak lensing come from DES Y3 (red) and KiDS-1000 (black), and the CMB backlight used for stacking the kSZ signal is from the ACT DR5 (orange) survey. We mask the full ACT footprint using the Planck $f_{\rm sky}=60\%$ galactic mask from Planck Collaboration et al. (2020a). We estimate that DES Y3, KiDS-1000, and ACT have an overlap area of 777 deg2, 409 deg2, and 5737 deg2 with BOSS respectively. Right: Redshift distributions for the galaxy surveys used in this analysis from Amon et al. (2023) (BOSS scaled by $10^{3}$). The BOSS CMASS survey is divided into 2 samples C1 (pink) and C2 (purple) by redshift. DES Y3 and KiDS-1000 survey bins are coloured to indicate which lens bin they correspond to, such that the sources are sufficiently behind the lenses.
@@ 2.1 Galaxy-Galaxy Lensing
The KiDS-1000 weak lensing data were taken with the 2.6 m VLT Survey Telescope (VST) (Capaccioli and Schipani, 2011; Grado et al., 2012) and span 777 $\rm deg^{2}$ of the sky (Giblin et al., 2021) with a number density of 6.22 $\rm arcmin^{-2}$.
The spectroscopic lenses from BOSS CMASS are subdivided into 2 lens redshift bins: C1 corresponding to $0.43<z<0.54$ and C2 corresponding to $0.54<z<0.70$. Each lensing survey (DES Y3 and KiDS-1000) providing background source galaxy shapes is also divided into 4 tomographic redshift bins, of which, we only use tomographic redshift bins sufficiently behind the foreground lens redshift bins (a GGL measurement of lens bin C1 uses both tomographic bins [3,4] while C2 uses only the [4] tomographic bin). The redshift distributions $n(z)$ of the lenses and sources are shown in the top-right panel of Figure 1. The tomographic bins used in DES Y3 and KiDS-1000 are calibrated using a tool from machine-learning known as the self-organizing map (SOM) (Buchs et al., 2019; Wright et al., 2020) which requires spectroscopy to accurately estimate photometric redshifts in each tomographic bin.
@@
Amon et al. (2023) create the DES Y3+KiDS-1000 combined data vector used in this analysis by taking the inverse-variance weighted average of GGL measurements made using DES Y3 and KiDS-1000 separately. They show that these measurements can be combined because the KiDS-BOSS and DES-BOSS on-sky footprints have no overlap (see Figure 1) and they are statistically consistent (see Section 5.1.5 of Amon et al. (2023) for more detail).
The error in the DES Y3 measurement is estimated using a Jackknife covariance. The lens galaxies and random points are split into 75 regions within the overlap area of 770 deg2 using the kmeans algorithm, resulting in each square region spanning 3 deg on a side or $\sim 65\;h^{-1}\rm Mpc$ (30 % larger than the largest angular scale measured). The KiDS-1000 error estimation is computed using the bootstrap method with square regions of 4 deg on each side equivalent to $\sim 80\;h^{-1}\;\rm Mpc$.
@@
The kSZ effect measurements used in this work were obtained by Schaan et al. (2021) by stacking the CMB temperature fluctuations mapped by ACT (in 98 GHz and 150 GHz frequencies) around BOSS CMASS galaxies (DR10), weighted by their peculiar velocities. The covariance matrix is constructed using the bootstrap method. [^2] Higher signal-to-noise kSZ measurements have recently been reported in (Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b) using ACT DR6 data; our method could be extended in the future to incorporate these data.
@@
The measurements span a range of aperture radii $\theta_{\rm d}$ from 1 to 6 arcminutes which probes the kSZ profile in the range of 0.5 to 4 virial radii in the CMASS galaxy sample.
These scales were chosen to probe baryonic feedback while minimizing covariance in the measurements. Schaan et al. (2021) stacked the measured temperatures $\mathcal{T}_{i}(\theta_{\rm d})$ around each galaxy in the sample. The stacking is done using a velocity weighted, inverse-variance weighted mean estimator given by
@@
where $\mathbf{v}$ is the velocity field, $f=d\ln\delta/d\ln a$ is the logarithmic linear growth rate, $\mathbf{\hat{n}}$ is the line-of-sight, $a$ is the scale factor, $H$ is the Hubble parameter, $b$ is the linear bias, and $\delta_{\rm g}$ is the galaxy overdensity.
Solving the linearized continuity equation allows for the reconstruction of the radial velocity field which can be assigned to each galaxy in the sample. The estimator is carefully chosen so that the radial direction of the velocity field does not cancel out thus leaving the kSZ signal intact. Lastly, the measurements of Schaan et al. (2021) perform a bootstrap resampling of the individual galaxies to estimate the covariance of the stacked kSZ profiles. Specifically, Schaan et al. (2021) repeatedly draw individual galaxies from the galaxy catalogue to generate a resampled galaxy catalogue of the same size. The kSZ profile is then measured on 10,000 realizations of the resampled galaxy catalogue which enable an estimation of the covariance matrix. The measurements at smaller aperture sizes have a dominant source of noise coming from the detector and foregrounds while the measurements at larger aperture sizes are dominated by large-scale CMB fluctuations.
@@ 3 Halo model
The first term refers to the 1-halo density profile which corresponds to density from the effective central halo, and the second term refers to the 2-halo density profile which corresponds to contributions in the density coming from the surrounding neighbouring halos (Asgari et al., 2023). We model the transition between the 1-halo and 2-halo terms as a sum rather than taking the maximum of the terms at any given scale. This simplistic treatment of the transition region should suffice for the current level of precision in our analysis, but future work that improves the precision of this model should explore the impact of this modelling choice. In the subsequent subsections below we will explain these 1-halo and 2-halo terms in more detail.
[…]
@@ 3.1 The One Halo Term
The 1-halo term of the density profile is the contribution to the density from the central halo itself. For this work we assume the host galaxy lies at the centre of a dark matter halo and that the gas is bound to the dark matter halo (i.e. no miscentering). In other words, we do not model the gas and dark matter around satellites and centrals with a Halo-Occupation Distribution (HOD) framework. Additionally, this model ignores the potential systematic effect of satellites or ‘miscentering’, where satellite galaxies offset the stacking procedures used on the central galaxies in both the kSZ effect measurements and the GGL measurements.
The impacts of ignoring satellites on the kSZ and GGL signals have been shown in (Hadzhiyska et al., 2023; McCarthy et al., 2024; Siegel et al., 2025; Bigwood et al., 2025). For CMASS galaxies Siegel et al. (2025) showed that ignoring satellites shifts halo mass estimates at the $\sim 10\%$ level, which is not significant enough to undermine the findings of this work. McCarthy et al. (2024) and Bigwood et al. (2025) showed with hydro simulations that for a CMASS-like sample in hydro simulations satellite fractions can be $\sim 10-20\%$ and have a scale dependent influence on the amplitude of both the kSZ and GGL signals at the $\sim 20\%$ level.
We also ignore an additional uncertainty on the reconstruction bias factor $r_{v}$ which is about $\sim 10\%$ for the reconstruction used in this analysis (Schaan et al., 2021). We note that modelling of more precise measurements should include this effect (Ried Guachalla et al., 2024; Hadzhiyska et al., 2024a; Ried Guachalla et al., 2025b).
The central goal of this work is to demonstrate the complementarity of measurements of the kSZ effect and GGL. This goal does not require a level of precision that warrants the accounting of these additional systematic effects. For this reason, it is suitable to make these modelling simplifications in this first iteration of the model framework. In Sections 4.2 and 4.3 we quantify how much these simplifications are expected to impact our parameter inference and by extension our analysis results. Future work designed for precise cosmological inference should account for these effects.
@@ 3.1.1 Dark Matter
to determine the concentration given a halo mass $c(M_{\rm halo})$ at $z=0.55$. By using $\Delta=200$, our definition of halo mass $M_{\rm halo}$ is the same as $M_{200}$. This concentration relation was determined from N-body simulations for halo masses $10^{11}-10^{15}\;\rm M_{\odot}$ and redshifts $0<z<2$ which matches our selected sample. We use a comoving $\bar{\rho}_{\rm m}$ because the radius $r$ is in comoving units. With this
@@
In this work, we don’t model the potential effects of assembly bias from the surrounding environment in our CMASS sample. This effect could bias our results at the few percent level (Yuan et al., 2021; Amon et al., 2023) and is a necessary improvement for future work. We also choose to ignore the back-reaction of baryonic feedback on the dark matter distribution: a second-order effect which is not necessary to model for our current level of precision (van Daalen et al., 2011a; Springel et al., 2018). Our model of the 1-halo term closely follows an approach from Cromer et al. (2022) which modelled baryons in WL of clusters. This approach differs from HOD based halo models commonly used to model GGL (Nishimichi et al., 2019; Miyatake et al., 2022; Zacharegkas et al., 2022; Dvornik et al., 2023), but it does allow for an explicit modelling of the baryon density profile which can be used to jointly analyse the kSZ measurements enabling constraints on baryon density parameters and dark matter parameters simultaneously.
@@
$x_{c}$ Core scale fraction $\mathcal{U}(0.0,1.0)$ >0.48>0.48 >0.46>0.46 >0.49>0.49
@@
$\beta$ Outer slope $\mathcal{U}(1,5)$ $2.47^{+0.34}_{-0.46}$ <3.51<3.51 $2.45^{+0.30}_{-0.23}$
@@
$A_{\rm 2h}$ Two halo term amplitude $\mathcal{U}(0.0,1.5)$ >0.631>0.631 $0.66\pm 0.16$ $0.64\pm 0.16$